For the function tan^(βˆ’1) π‘₯Β  find the infinite Taylor series at π‘Ž = 0, the radius of convergence, range of convergence, derivative, integral and the product with itself. Β 


1
Expert's answer
2021-12-21T10:46:21-0500

"\text{Recall well-known series},
\\

\frac{1}{1-x} =1+x+x^{2} +\cdots =\sum _{n=0}^{\infty }x^{n} , \left|x\right|<1
\\

\begin{array}{l} {\Rightarrow \tan ^{-1} \left(x\right)=\int \frac{1}{1+x^{2} } dx =\int \frac{1}{1-\left(-x^{2} \right)} dx } \\ {=\int \sum _{n=0}^{\infty }\left(-x^{2} \right)^{n} dx ,\left|x^{2} \right|<1\Rightarrow \left|x\right|^{2} <1\Rightarrow \left|x\right|<1} \\ {=\int \sum _{n=0}^{\infty }\left(-1\right)^{n} x^{2n} dx ,\left|x\right|<1} \\ {=\sum _{n=0}^{\infty }\frac{\left(-1\right)^{n} x^{2n+1} }{2n+1} ,\left|x\right|<1} \end{array}
\\
\Rightarrow R=1
\\

\text{When } x=-1,\sum _{n=0}^{\infty }\frac{\left(-1\right)^{n} x^{2n+1} }{2n+1} =\sum _{n=0}^{\infty }\frac{\left(-1\right)^{n+1} }{2n+1} \text{which converges by Alternating series test as } \frac{1}{2n+1} \to 0 \text{ and it also decreasing}
\\

\text{When } x=1,\sum _{n=0}^{\infty }\frac{\left(-1\right)^{n} x^{2n+1} }{2n+1} =\sum _{n=0}^{\infty }\frac{\left(-1\right)^{n} }{2n+1} \text{which converges by Alternating series test as } \frac{1}{2n+1} \to 0 \text{ and it also decreasing}
\\

\text{Hence, the Radius of convergence } R=1
\\

\text{The Range of convergence } I=[-1,1]
\\
\text{The derivative } \frac{d}{dx} \left(\tan ^{-1} x\right)=\frac{1}{1+x^{2} } 
\\

\text{ To obtain the Integral, we shall use Integration by part, }
\\

\text{Let } J=\int \tan ^{-1} \left(x\right)dx , u=\tan ^{-1} \left(x\right) \quad dv=dx
\\
\hspace{98pt} du=\frac{1}{1+x^{2} } dx \qquad v=x
\\
\Rightarrow J=\int \tan ^{-1} \left(x\right)dx =x\tan ^{-1} \left(x\right)-\frac{1}{2} \int \frac{2x}{1+x^{2} } dx 
\\
\hspace{95pt}=x\tan ^{-1} \left(x\right)-\frac{\ln \left(x^{2} +1\right)}{2} +C
\\
\text{The Product}
\\
{\left( {{{\tan }^{ - 1}}x} \right)^2} = {\left( {x - \frac{1}{3}{x^3} + \frac{1}{5}{x^5} + \cdots } \right)^2} = {x^2} - \frac{2}{3}{x^4} + \frac{{23}}{{45}}{x^6} - \cdots"